OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7. y = 4x + 10 y = one fourthx + 5 halves y = -4x - 6 y = -one fourthx + 3 halves

OpenStudy (anonymous):

slope of line parallel to y = 4x + 7. is m=4 let our equation of line is y=mx+c y=4x+c as it passes thrupoint (-2, 2) 2=-8+c c=10 line is y=4x+10